# HeadLighter: Disentangling Illumination in Generative 3D Gaussian Heads via Lightstage Captures

### v1

@@ 3.4 Disentanglement
Our goal is to disentangle intrinsic head appearance and extrinsic illumination in a generative 3D Gaussian head model, supervised by high-fidelity lightstage captures. To achieve stable training, we devise a three-stage training strategy: (1) The first stage trains the base branch $G_{base}$, adapting a pretrained entangled generator to light-invariant domain. (2) Based on the guidance of the base branch, the second stage trains the relight branch $G_{relit}$, learning the interaction between incident light and head surfaces using the captured lightstage data. (3) The final stage jointly fine-tunes the entire model to harmonize the predictions from both branches and enhance overall realism.
3.4.1 Stage 1: Base Branch Adaptation
We initialize $G_{base}$ by adapting a pretrained head generative model, GGHead [gghead]. GGHead does not decouple intrinsic appearance and scene illumination, and its color output first-order SH coefficients per RGB channel represents a view-dependent radiance that entangles both factors. We repurpose the pretrained generator to produce a view-independent RGB albedo map, $\rho_{k}$, which represents the intrinsic appearance; while keeping the generator’s internal feature representation unchanged.
The modified generator is first finetuned on the FFHQ dataset using an adversarial loss, which enforces a view-independent radiance, disentangling illumination from the intrinsic appearance at the base level, while preserving generalization ability. Subsequently, it is finetuned for 10k iterations using pseudo albedo maps derived from [nerffacelighting] by adversarial loss. This stage adapts the generator to the domain of albedo (light-invariant), while the short training duration prevents identity collapse.
3.4.2 Stage 2: Relight Priors from LightStage Captures
@@
Stage 2 focuses on training the relight branch $G_{\text{relit}}$, while keeping the base generator $G_{\text{base}}$ frozen. We first implement multi-view GAN inversion on $G_{base}$ to obtain geometry and albedo of lightstage captures. We then optimize $G_{relit}$ and latent code per subject to predict relighting-related attributes $(\mathbf{n}_{k},\sigma_{k},\mathbf{T}_{d,k})$ that match the lightstage observations, where the multi-scale features output by $G_{base}$ are injected into $G_{\text{relit}}$ as a guidance. A novel normal distillation and normalization strategy is proposed to enable high-quality normal prediction. We further utilize regularization terms to constrain the generated attributes to be plausible.
@@
$\mathcal{L}_{inv}=\mathcal{L}_{2}+\mathcal{L}_{perc}+\mathcal{L}_{tv}.$ (5)
Notably, unlike methods such as PTI [roich2022pivotal] that fine-tune the generator weights per subject, our approach keeps the base generator $G_{\text{base}}$ fixed throughout inversion to preserve consistent feature representations across different identities in the latent space.
@@
$\mathbf{F}^{l}_{\text{relit}}=\mathbf{F}^{\prime l}_{\text{relit}}+\Delta^{l},\quad\Delta^{l}=\Phi^{l}\big(\mathbf{F}^{l}_{\text{base}},\mathbf{F}^{\prime l}_{\text{relit}}\big).$ (6)
@@
To ensure the correctness and smoothness of the predicted normals, we introduce a self-distillation loss and 2D smoothness term. Based on the inverted geometry represented by gaussian point clouds, we pre-extract meshes via poisson reconstruction and cleanup the noise hair region by projecting onto 2D skin-parsing [face-parsing] masks. Let $\hat{\mathbf{n}}^{k}_{\text{mesh}}$ be the closest mesh triangle normal of the gaussian splat $g_{k}$. We minimize the angular deviation between normalized normal prediction $\mathbf{n}_{k}$ and $\hat{\mathbf{n}}_{k}^{\text{mesh}}$ via cosine similarity:
$\mathcal{L}_{normal}^{D}=\frac{1}{N}\sum_{k=1}^{N}(1-\mathbf{n}_{k}^{\top}\hat{\mathbf{n}}_{k}^{\text{mesh}}).$ (7)
@@
$\mathcal{L}_{normal}^{TV}=\left\|\nabla_{x}I_{\text{norm}}\right\|_{1}+\left\|\nabla_{y}I_{\text{norm}}\right\|_{1},$ (8)
@@
$\mathcal{L}_{shading}=\mathcal{L}_{\text{img}}+\mathcal{L}_{perc}+\mathcal{L}_{tv}^{I}.$ (9)
@@
$\mathcal{L}_{\text{prt}}=\frac{1}{N}\sum_{k=1}^{N}\|\mathbf{T}_{k}-\bar{\mathbf{T}}_{k}\|_{2}^{2},\quad\bar{\mathbf{T}}_{k}=\frac{1}{3}\sum_{c\in\{r,g,b\}}\mathbf{T}_{k}^{(c)}.$ (10)
3.4.3 Stage3: Joint Training
In this final stage, we perform end-to-end training of the full model to further enhance the diversity and realism of generated head images under arbitrary light conditions. We retain the loss functions introduced in Stage 2 to ensure physically plausible relighting outputs from $G_{\text{relit}}$, while additionally incorporating an adversarial loss with a discriminator trained on the FFHQ dataset, to promote high-fidelity and diverse image generation.
In this stage, for adversarial training, we sample random lighting conditions, random base latent code, and random relight latent code, to synthesize albedo, geometry, and relighting-related attributes of a relightable 3DGS head using $G_{base}$ and $G_{relit}$. Since in Stage 2, the input to $G_{relit}$ is $\mathbf{w}^{+}$ code obtained from multi-view inversion. To support random sampling of relight latent code, we first train the mapping network $M_{\text{lit}}$, which maps a randomly sampled latent code $z_{lit}\in\mathbb{R}^{128}$ into the $\mathcal{W}^{+}$ space of $G_{relit}$, using the optimized $\mathbf{w}^{+}_{\text{lit}}$ obtained from Stage 2 as training data. To support random sampling of various lighting conditions, we employ an off-the-shelf light estimation method [zhou2019deep] to estimate SH coefficients for FFHQ images.
During each training iteration, we sample random lighting conditions together with random $\mathbf{z}^{+}$ (from $G_{\text{base}}$) and $\mathbf{z}^{+}_{\text{lit}}$ (from $M_{\text{lit}}$) to synthesize a relighted head image. The generated image is then evaluated by a discriminator, which guides the generator toward producing more photo-realistic and diverse outputs. This stage effectively bridges the gap between lightstage captures and in-the-wild images, allowing the model to explore a broader manifold of plausible head appearances under varying illumination.
4 Experiments
Figure 3: 3D head generation and lighting control. EG3D and GGHEAD do not support explicit lighting control. NFL and GSHR support lighting editing but struggle to handle extreme input lighting (strong, side light), while our method can effectively interact with such challenging conditions. Lighting conditions are provided at the lower right corner.
Figure 4: Relighting under complex environment maps. NFL assumes white illumination and fails to handle colored lighting. GSHR lacks real-world OLAT supervision and therefore cannot produce physically plausible responses under complex lighting. In contrast, our method, trained with multi-view OLAT captures, accurately renders realistic shading and reflections under diverse and challenging illumination conditions.
@@ 4 Experiments
Ours $\times$ 4.68 201
Table 1: Comparison of generative quality and rendering speed. SR denotes super resolution modules and Relit denotes whether to support explicit lighting control. Rendering speed is tested on Nvidia A6000 GPU.
4.1 Implementation Details.
Lightstage Dataset. We perform calibration to obtain the world coordinates of each point light and camera. For the captured multi-view images, we use open-source tools to detect 2D landmarks [bulat2017far] and human segmentation [zheng2024bilateral] to the ”Uniform” frame. All the captured frames are aligned and cropped into 512 resolution. The head pose for each frame is solved using multi-view 2D landmarks and calibrated camera extrinsics, constrained by a 3DMM prior [paysan20093d]. $56$ participants (31 male, 25 female) perform three expressions (neutral, smile, and eyes closed), each recorded under $95$ distinct lighting configurations. We select $46$ subjects for training and $10$ for testing.
Architecture and Training. $G_{\text{relit}}$ shares the same overall architecture and number of layers as the $G_{\text{base}}$. However, since $G_{\text{base}}$ models complex geometric and textural details that go far beyond face reflectance alone, we design $G_{\text{relit}}$ with reduced capacity: the convolutional layers in $G_{\text{relit}}$ operate on features of up to 256 dimensions, compared to up to 512 dimensions in $G_{\text{base}}$. Correspondingly, we use a 128-dimensional latent code $\mathbf{z}_{\text{lit}}$ to represent face material properties (e.g., albedo, normals, roughness), while geometry and intrinsic appearance are encoded by a 512-dimensional latent code $\mathbf{z}_{\text{base}}$. During Stage 2, $G_{\text{relit}}$ is optimized with Adam using a learning rate of $5\times 10^{-4}$, while the lighting latent code $\mathbf{w}_{\text{lit}}$ is updated with a higher learning rate of $1\times 10^{-3}$. In Stage 3, both are trained with a reduced learning rate of $1\times 10^{-4}$. Additional details are provided in the supplementary material.
@@
Table 2: Evaluation of spatial consistency of generated 3D heads. We render multi-view images from the generated 3D heads and synthesis novel view using NeuS2 [wang2023neus2]; higher PSNR, SSIM, and lower LPIPS indicate better spatial consistency.
Figure 5: Free view relighting on captured lightstage data. We show the lighting condition at the lower right corner.
4.2 3D Generation.
We evaluate generation quality against multiple state-of-the-art 3D-aware head synthesis methods from both qualitative and quantitative perspectives. First, we assess generation diversity by generating 50K images and computing FID against the FFHQ dataset [kazemi2014one]. Second, we measure rendering speed on one NVIDIA A6000 GPU. Third, we evaluate multi-view spatial consistency: for each random latent code, we render 30 viewpoints across $45\degree$ to $45\degree$, hold out the frontal view as ground truth to calculate metrics, and use the remaining views to reconstruct the geometry by NeuS2 [wang2023neus2].
As shown in Fig 3, we compare against four state-of-the-art methods: EG3D [eg3d] and NFL [nerffacelighting], which rely on neural volume rendering, and GGHead [gghead], GSHR [lv2025gsheadrelight], which are built upon 3D Gaussian splatting. Among them, EG3D and GGHead do not support explicit illumination control. NFL and GSHR support lighting control, but due to the lack of real-world data, they struggle to respond accurately to complex or extreme lighting conditions. We show the FID score and rendering FPS in Table 1. Methods based on 3D Gaussian achieve significantly higher rendering speeds than volume-rendering counterparts. Our method is slightly slower than GGHead due to the inclusion of a physically-based shading model but maintains the real-time rendering performance. We evaluate multi-view consistency using a wide viewport range (±45°). As shown in Table 3, our method achieves the highest multi-view consistency. Our model is trained on multi-view lightstage face images, which encourages consistency across viewpoints.
4.3 Light Control.
As shown in Figure 4, we compare the lighting control capabilities of different methods. Due to the lack of real-world supervision, NFL achieves disentanglement under a white-light assumption combined with style-mixing, and thus cannot handle colored incident lighting. GSHR is an unsupervised disentanglement method based on 3D Gaussians and it also fails to handle complex or extreme lighting conditions. Our method can handle complex, colored incident lighting with high photorealism. We provide additional comparisons with other methods on lighting control in the supplementary material.
@@
Ours 26.0923 0.8514 0.0988
Table 3: Quantitative comparisons of free view relighting on lightstage captures.
4.4 Portrait Relighting.
We validate the model’s ability to disentangle illumination and appearance using real-captured data. We extract image pairs from lightstage captures and perform free-view relighting using our method and baselines. The first row in the figure shows relighting from uniform illumination to a side-light condition, while the second row demonstrates delighting from non-uniform, dim illumination back to uniform lighting. Compared to baseline methods, our approach responds more accurately to the input lighting and better preserves the subject’s geometry and appearance. We also provide quantitative results in Tab 3
Figure 6: Normal generation with different methods.
Figure 7: Face Material Editing.
4.5 Normal Distillation.
We show the normal calculation results of different methods in Fig 6. Directly computing normals from neighboring points in gaussian splats introduces significant noise, making it unsuitable for photorealistic rendering. Reconstructing a mesh from the Gaussian point cloud and then extracting normals is computationally expensive to the training process and often produces noticeable artifacts in fine regions such as hair. To address this, we directly predict normals using the $G_{relit}$ branch, supervised by normals derived from a pre-extracted mesh combined with a 2D total variation loss, yielding smooth and accurate surface normals.
4.6 Face Material Editing.
Thanks to the disentangled representation of intrinsic face materials in our framework, we enable explicit editing of facial appearance properties independent of geometry and illumination. Figure 7 shows a set of results demonstrating control over facial oiliness by adjusting roughness.
5 Conclusion.
@@
Limitations.
The model is static and does not support the control of facial expressions. Furthermore, the number of captured identities and lighting conditions remains limited, which restricts the model’s ability to learn a more comprehensive prior over diverse facial material properties. Future work could address these by incorporating video captures and expanding the dataset scale.
[…]
Supplementary Material
6 Implementation Details
Figure 8: An overview of the lightstage captures.
Dataset. We provide a summarized overview of the LightStage dataset in Fig 8. The dataset encompasses diverse subjects, lighting conditions, facial expressions, and viewing angles. The lighting conditions are categorized into three types: uniform, directional, and random. The facial expressions include neutral, smiling, and closed eyes.
Training. We provide hyper parameters and configurations of training in this section.
  • Stage1: We retain the majority of the pretrained GGHead model’s architecture and parameters, and modify only the final layer to output RGB values instead of the original spherical harmonic coefficients. We fine-tune this modified model for 80,000 iterations on the FFHQ dataset using an adversarial loss. We then generate 10,000 pseudo-albedo images using NFL [nerffacelighting]. To prevent identity collapse, we use a large truncated $\psi$ value, which controls the diversity of sampled latent codes by constraining the sampling region in the latent space. We fine-tune the model for an additional 10,000 iterations on these pseudo-albedo images with an adversarial loss, removing the lighting component from the pretrained GGHead model while maintaining the diversity of the generated identities.
  • Stage2-Inversion: We achieve multi-view inversion under the supervision of multiple loss terms. We set the weight of the $\mathcal{L}_{2}$ to 0.1, the weight of the $\mathcal{L}_{perc}$ to 1, and the weight of $\mathcal{L}_{tv}$ to 0.01. We compute the perceptual loss only using frontal images. The inversion process requires 500 iterations, and the inversion result of the previous fully lit frame serves as the initial value for the next frame to accelerate convergence.
  • Stage2-Network: We employ a lightweight network including 2 layers and Leaky ReLU activation to enable multi-scale features of $G_{base}$ to guide the relight branch. We design $G_{\text{relit}}$ with a reduced feature dimensionality: its convolutional layers operate on 256-dimensional features, compared to 512 in $G_{\text{base}}$. Finally, the relight branch requires approximately 90 MB memory storage.
  • Stage2-Training: We set weights of $\mathcal{L}_{normal}^{D}$, $\mathcal{L}_{normal}^{TV}$ to be 1 and 10 respectively. The weights of $\mathcal{L}_{img}$, $\mathcal{L}_{perc}$ and $\mathcal{L}_{tv}^{I}$ are $1$. And weight of $\mathcal{L}_{prt}$ is set to $0.5$.
  • Stage3-Training: Firstly, we train mapping network $M_{lit}$ for 10k iterations. Then we train the two branches using losses mentioned in Stage 2 and additional adversarial loss. The parameters of discriminator are fixed.
Figure 9: Comparisons with LumiGAN [deng2024lumigan].
7 More Comparisons
We also provide comparisons with FaceLit [ranjan2023facelit] in Fig 10 and LumiGAN [deng2024lumigan] in Fig 9. Both FaceLit and LumiGAN are unsupervised disentanglement methods based on volume rendering, which struggle with extreme lighting conditions and often produce unrealistic lighting artifacts. Moreover, FaceLit relies on a white-light assumption and therefore cannot handle colored lighting. In contrast, our method supports diverse lighting inputs, including color environment maps, point lights, and spherical harmonic coefficients. It also produces natural lighting and shadow effects even under strong side lighting conditions.
Figure 10: Comparisons with FaceLit [ranjan2023facelit].
8 Generation Results.
Fig 11 shows additional identities generated under various lighting conditions. Our method produces diverse facial materials and supports relighting under low-light environments and multiple environmental maps.
Figure 11: More Generation Results. We show the corresponding environment maps at the right down corner.
9 Ethical Considerations
The generation of artificial portrait using our method poses risks, including the spread of false information, and erosion of trust in media credibility. These issues could have profound societal implications. Addressing this challenge requires developing reliable techniques to identify and verify authentic content.

### v2

@@ 3.4. Disentanglement
Our goal is to disentangle intrinsic head appearance and extrinsic illumination in a generative 3D Gaussian head model, supervised by high-fidelity lightstage captures. To achieve stable training, we devise a three-stage training strategy: (1) The first stage trains the base branch $G_{base}$, adapting a pretrained entangled generator to light-invariant domain. (2) Based on the guidance of the base branch, the second stage trains the relight branch $G_{relit}$, learning the interaction between incident light and head surfaces using the captured lightstage data. (3) To generalize our model to arbitrary illumination beyond the limited white-light OLAT data, we apply adversarial loss on images rendered with randomly sampled lighting.
3.4.1. Stage 1: Base Branch Adaptation
We initialize $G_{base}$ by adapting a pretrained head generative model, GGHead (Kirschstein et al., 2024) to produce $\mathcal{A}_{\text{geo}}$ and view-independent albedo $\rho_{k}$, while maintaining its diversity of high-quality generation. GGHead does not decouple intrinsic appearance and scene illumination, and its color output first-order SH coefficients per RGB channel represents a view-dependent radiance that entangles both factors. We firstly modify GGHead to output 3-channel RGB colors instead of the original 12-channel SH coefficients. The modified generator is first finetuned on the FFHQ dataset (with segmentation masks derived from SAM (Kirillov et al., 2023) and (Zheng et al., 2024)) using an adversarial loss, which enforces a view-independent radiance, discarding anisotropic shading components at the base level. Subsequently, it is finetuned for 20k iterations using pseudo albedo images generated by (Jiang et al., 2023) via adversarial loss. Instead of updating all layers, we restrict updates to the intermediate layers (specifically, layer 5 and 6 in the GGHEAD backbone). These intermediate layers are known to encode mid-level appearance features such as shading and illumination (Karras et al., 2019). This targeted adaptation steers the generator toward the albedo domain while preserving high-level identity structure and face details, and the limited training duration further mitigates the risk of collapsing to the limited pseudo albedo images.
3.4.2. Stage 2: Relight Priors from LightStage Captures
@@
Stage 2 focuses on training the relight branch $G_{\text{relit}}$, while keeping the base generator $G_{\text{base}}$ frozen. We first implement multi-view GAN inversion on $G_{base}$ to obtain geometry and albedo of lightstage captures. We then optimize $G_{relit}$ and latent code per subject to predict relighting-related attributes $(\mathbf{n}_{k},\sigma_{k},\mathbf{T}_{d,k})$ that match the lightstage observations, where the multi-scale features output by $G_{base}$ are injected into $G_{\text{relit}}$ as a guidance. A novel normal distillation strategy is proposed to predict high-quality normal. We further utilize several regularization terms to constrain the generated attributes to be plausible.
@@
(6) $\mathcal{L}_{inv}=\mathcal{L}_{2}+\mathcal{L}_{perc}+\mathcal{L}_{tv}.$
Notably, unlike methods such as PTI (Roich et al., 2022) that fine-tune the generator weights per subject, our approach keeps the base generator $G_{\text{base}}$ fixed throughout inversion to preserve consistent feature representations across different identities in the latent space.
@@
(7) $\mathbf{F}^{l}_{\text{relit}}=\mathbf{F}^{\prime l}_{\text{relit}}+\Delta^{l},\quad\Delta^{l}=\Phi^{l}\big(\mathbf{F}^{l}_{\text{base}},\mathbf{F}^{\prime l}_{\text{relit}}\big).$
@@
To ensure the correctness and smoothness of the predicted normals, we introduce a self-distillation loss and 2D smoothness term. Based on the inverted geometry represented by gaussian point clouds, we pre-extract meshes via poisson reconstruction and cleanup the noise hair region by projecting onto 2D skin-parsing (Yakhyokhuja, 2024) masks. Let $\hat{\mathbf{n}}^{k}_{\text{mesh}}$ be the closest mesh triangle normal of the gaussian splat $g_{k}$. We minimize the angular deviation between normalized normal prediction $\mathbf{n}_{k}$ and $\hat{\mathbf{n}}_{k}^{\text{mesh}}$ via cosine similarity:
(8) $\mathcal{L}_{normal}^{D}=\frac{1}{N}\sum_{k=1}^{N}(1-\mathbf{n}_{k}^{\top}\hat{\mathbf{n}}_{k}^{\text{mesh}}).$
@@
(9) $\mathcal{L}_{normal}^{TV}=\left\|\nabla_{x}I_{\text{norm}}\right\|_{1}+\left\|\nabla_{y}I_{\text{norm}}\right\|_{1},$
@@
(10) $\mathcal{L}_{shading}=\mathcal{L}_{\text{img}}+\mathcal{L}_{perc}+\mathcal{L}_{tv}^{I}.$
@@
(11) $\mathcal{L}_{\text{prt}}=\frac{1}{N}\sum_{k=1}^{N}\|\mathbf{T}_{k}-\bar{\mathbf{T}}_{k}\|_{2}^{2},\quad\bar{\mathbf{T}}_{k}=\frac{1}{3}\sum_{c\in\{r,g,b\}}\mathbf{T}_{k}^{(c)}.$
Figure 3. Relighting under complex environment maps. LuminGAN (Deng et al., 2024) and NFL assumes white illumination and fails to handle colored lighting. GSHR lacks real-world OLAT supervision and therefore cannot produce correct responses under non-uniform colored lighting. In contrast, our method accurately reproduces spatially varying lighting effects.
3.4.3. Stage3: Adversial Training
In this final stage, we perform adversial training to ensure robust photorealism under arbitrary illumination. Given that our dataset consists solely of limited white-light OLAT captures, we employ adversarial supervision on images generated with randomly sampled illumination. This strategy expands the training distribution, ensuring the model generalizes well to diverse illumination environments.
First, to support random sampling of the relight latent code, we initialize mapping network $M_{\text{lit}}$ that maps a randomly sampled code $z_{lit}\in\mathbb{R}^{128}$ into the $\mathcal{W}^{+}$ space of $G_{relit}$, utilizing the optimized $\mathbf{w}^{+}_{\text{lit}}$ from Stage 2 as training data. Subsequently, we freeze the parameters of the base branch $G_{base}$, exclusively optimizing the synthesis network of $G_{relit}$. During each training iteration, we simulate diverse illumination conditions to broaden the lighting manifold beyond the training set. We pre-define a set of candidate point lights uniformly distributed on the hemisphere frontal to the face. A random lighting condition $\mathcal{L}_{rand}$ is constructed by randomly sampling a subset of candidate lights and assigning random RGB intensities to each source. Simultaneously, we sample random latent codes $\mathbf{z}_{base}$ for the geometry representation and random $\mathbf{z}_{lit}$ for the materials. The relight branch $G_{relit}$ synthesizes the relighted head image using these random inputs, which is then supervised by an adversarial loss using pretrained discriminator of GGHead. This supervision guides the generator to produce high-fidelity, photorealistic textures that remain stable under varying, colored synthetic lighting. Moreover, the losses from Stage 2 are retained in this final stage to serve as a regularizer to prevent the relight branch from drifting. To ensure stable fine-tuning, we employ a reduced learning rate of $1\times 10^{-6}$.
4. Experiments
@@
Ours $\times$ 4.78 201
Table 1. Comparison of generative quality and rendering speed. SR denotes super resolution modules and Relit denotes whether to support explicit lighting control. Rendering speed is tested on Nvidia A6000 GPU.
4.1. Implementation Details.
Lightstage Dataset. We first perform calibration to obtain the world coordinates of each point light and camera. For the captured multi-view images, we use open-source tools to detect 2D landmarks (Bulat and Tzimiropoulos, 2017) and human segmentation (Zheng et al., 2024) for the ”Uniform” frame. All the frames are aligned and cropped into 512 resolution. The head pose for each frame is solved using multi-view 2D landmarks and calibrated camera extrinsics, constrained by a 3DMM prior (Paysan et al., 2009). $56$ participants (31 male, 25 female) perform three expressions (neutral, smile, and eyes closed), each recorded under $95$ distinct lighting configurations. We select $46$ subjects for training and $10$ for testing.
Architecture and Training. $G_{\text{relit}}$ shares the same overall architecture and number of layers as the $G_{\text{base}}$. However, since $G_{\text{base}}$ models complex geometric and appearance details that go far beyond head reflectance alone, we design $G_{\text{relit}}$ with reduced capacity: the convolutional layers in $G_{\text{relit}}$ operate on features of up to 256 dimensions, compared to up to 512 dimensions in $G_{\text{base}}$. Correspondingly, we use a 128-dimensional latent code $\mathbf{z}_{\text{lit}}$ to represent head material properties, while geometry and intrinsic appearance are encoded by a 512-dimensional latent code $\mathbf{z}_{\text{base}}$. During Stage 2, $G_{\text{relit}}$ is optimized with Adam using a learning rate of $5\times 10^{-4}$, while the lighting latent code $\mathbf{w}_{\text{lit}}$ is updated with a higher learning rate of $1\times 10^{-3}$. Additional details are provided in the supplementary material.
4.2. 3D Generation.
We evaluate generation quality against multiple state-of-the-art 3D-aware head synthesis methods. First, we assess generation diversity by generating 50K images and computing FID against the FFHQ dataset (Kazemi and Sullivan, 2014). Images are rendered under random lighting conditions synthesized as described in Stage 3. Second, we measure rendering speed on a NVIDIA A6000 GPU. Third, we evaluate multi-view spatial consistency: for a random latent code, we render 30 viewpoints across $45\degree$ to $45\degree$, hold out the frontal view as ground truth to calculate metrics, and use the remaining views to reconstruct the geometry by NeuS2 (Wang et al., 2023).
We compare with four baselines: EG3D (Chan et al., 2022) and NFL (Jiang et al., 2023), which rely on neural volume rendering, and GGHead (Kirschstein et al., 2024), GSHR (Lv et al., 2025), which are built upon 3D Gaussian splatting. Among them, EG3D and GGHead do not support explicit illumination control while NFL and GSHR support lighting control. We show the FID score and rendering FPS in Table 1. Methods based on 3D Gaussian achieve significantly higher rendering speeds than volume-rendering counterparts. Our method is slightly slower than GGHead due to the physically-based shading model but maintains the real-time rendering performance. We evaluate multi-view consistency using a wide viewport range (±45°). Our method achieves plausible disentanglement, which eliminates the inconsistent anisotropic artifacts, thus produces better multi-view consistency, as shown in Table 3.
@@
Table 2. Evaluation of spatial consistency of generated 3D heads. We render multi-view images from the generated 3D heads and synthesis novel view using NeuS2 (Wang et al., 2023); higher PSNR, SSIM, and lower LPIPS indicate better spatial consistency.
4.3. Light Control.
We compare the rendering performance of various generative methods under complex and extreme illumination conditions. The baselines include LumiGAN (Deng et al., 2024), NFL (Jiang et al., 2023), and GSHR (Lv et al., 2025). While LumiGAN and NFL are restricted by a white-light assumption , both GSHR and our method support colored illumination. Since LumiGAN’s code is unavailable, its results are sourced from the original paper. Figure 3 illustrates the results under non-uniform colored environment maps. Our generated 3D heads accurately respond to spatially-varying environmental lighting. Furthermore, Figure 4 demonstrates relighting results under extreme conditions, such as side-lighting or low-light environments. Lacking supervision from real-world data, NFL and GSHR fail to handle these extreme cases, either producing significant artifacts or failing to respond. In contrast, our method produces high-fidelity shadow and specular highlights.
4.4. Portrait Relighting.
Our method, along with NFL and GSHR, enable generating 3D human heads and support portrait illumination and viewport editing via 3D GAN inversion. To validate our model’s ability to disentangle illumination and appearance, we extract image pairs from LightStage captures and perform free-view relighting using our method and baselines. We also compare against Lite2Relight(Rao et al., 2024), a recent 3D GAN based portrait editing method. For 10 test subjects, we randomly sample two types of OLAT image pairs: one from uniform illumination to side lighting (as shown in the first row of Fig 5), and the other from side lighting to uniform illumination (shown in the second row of Fig 5), with 10 random lighting combinations for each type. Our method demonstrates superior robustness in handling low-light environments and extreme side-lighting. We also provide quantitative results in Tab 3.
@@
Lite2Relight 20.8874 0.8073 0.1278
Ours 27.0550 0.8614 0.1001
Table 3. Quantitative comparisons of free view relighting.
4.5. Ablation Study.
Base Branch Adaption. We fine-tune GGHead with NFL-generated pseudo-albedo images to remove illumination, optimizing only intermediate layers to avoid identity collapse. We compare results before and after optimization using the same latent code, as shown in Fig 6. From left to right: the original GGHead output, optimization of layer 6, optimization of two intermediate layers 5 and 6, and full-model optimization. Full-model optimization significantly alters facial geometry, biasing it toward the average face. Optimizing only the 6th layer fails to fully remove illumination artifacts. Our method effectively removes illumination while preserving coarse face structure and fine-grained details.
Normal Distillation. We show the normal calculation results of different methods in Fig 9. Directly computing normals from neighboring points in gaussian splats introduces significant noise, making it unsuitable for photorealistic rendering. Reconstructing a mesh from the Gaussian point cloud and then extracting normals is computationally expensive to the training process and often produces noticeable artifacts in fine regions such as hair. To address this, we directly predict normals using the $G_{relit}$ branch, supervised by normals derived from a pre-extracted mesh combined with a 2D total variation loss, yielding smooth and accurate surface normals.
Random Light and Adversarial Training. We evaluate the impact of random illumination and adversarial training by comparing rendering results under novel, colored lighting conditions not present in the white-light OLAT training set. As shown in Fig 8, without this stage, the model exhibits noticeable artifacts and color inconsistencies when subjected to non-uniform illumination. By incorporating adversarial training with randomly sampled lighting, our method ensures well generalization to diverse lighting, producing stable and realistic relighting results under challenging, unseen conditions.
Explicit Control on Roughness. Our method generates disentangled and physically-based face material properties. To evaluate the quality of disentanglement, we perform direct attribute editing on the roughness component. As illustrated in Figure 7, decreasing the roughness value leads to more pronounced specular highlights while preserving other appearance attributes, demonstrating effective decoupling of the material properties.
5. Conclusion.